OpenStudy (anonymous):

4x - y = 8 -8x + 2y = -5

OpenStudy (anonymous):

system of equations

OpenStudy (anonymous):

Just like in the previous question. Isolate any variable in either equation. It would be easier to isolate the y variable in the first equation. 4x - y = 8 => when you solve it for y, you get => y=4x-8 Now plug in "y=4x-8" into the second equation -8x + 2y = -5 Plug in "4x-8" where there is currently a Y. -8x+2(4x-8)=-5. Now simpllify this equation.

OpenStudy (anonymous):

But the problem you may encounter is that you won't get a solution. This is because these two lines are parallel, which means that they don't intersect. There would be no solution.